OpenStudy (anonymous):

What are the x-intercept(s) of the graph of y − x2 = 11x + 30? (−5, 0) and (6, 0) (5, 0) and (−6, 0) (−5, 0) and (−6, 0) (5, 0) and (6, 0)

OpenStudy (campbell_st):

you have a curve y = x^2 + 11x + 30 find the factors of 30 that add to 11

OpenStudy (anonymous):

(5, 0) and (−6, 0)? @campbell_st

OpenStudy (anonymous):

what do you mean I have a curve?

OpenStudy (campbell_st):

then solve x^2 + 11x + 30 = 0 (x +5)(x + 6) = 0 then x = -5, -6 these are the intercepts the graph is a curve... specifically... a parabola...

OpenStudy (anonymous):

Oh, so it would be (−5, 0) and (−6, 0)

OpenStudy (campbell_st):

yes